OpenStudy (anonymous):

what is the oxidation number of Ni(s) after the reaction: Fe (s) + NiCl2 (aq)  2 Ni (s) + Fe Cl2 (aq)?

OpenStudy (anonymous):

the equation is not balanced

sam (.sam.):

\[\begin{array}{c} \text{Fe} \\\end{array}\text{ + }\begin{array}{c} \text{Ni}\text{Cl}_2 \\\end{array}\text{ $\longrightarrow $ }\begin{array}{c} \text{Ni} \\\end{array}\text{ + }\begin{array}{c} \text{Fe}\text{Cl}_2 \\\end{array}\]

sam (.sam.):

From NiCl2 the oxidation number of Ni is +2, it then turns to neutral atom 0, its reduced to 0